A variable contains five categories. It is expected that data are uniformly distributed across these five categories. To test this, a sample of observed data is gathered on this variable resulting in frequencies of 27, 30, 29, 21, and 24. Using α = .01, the degrees of freedom for this test are ___.

5

4

3

2

1
Answer #1

Number of categories is 5 = n

Degrees fo freedom is n-1=5-1=4

chi square distribution is applied